# Quantum memory assisted observable estimation

### v3

@@ Appendix C Cost of implementing CPS
If we want the CPS method for observable estimation to be as generally applicable to the results of quantum algorithms, we need to satisfy the following two requirements. First, the implemented controlled reflection and controlled Pauli operation required for $c\hat{U}_{P_{j}}$ are not more error-prone than the state preparation unitary $V$. Second, we need to take into account the effect of applying the state preparation unitaries sequentially for the CPS method, as opposed to a shot-noise limited multi-shot estimation scheme where the system is measured after each application of $V$. The contrast between the sequential and multi-shot estimation protocols with respect to the accumulation of errors will be discussed in Appendix D below. In this section, we will focus on the implementation of controlled reflection and Pauli operations.
@@
Ideally, the implementation of the controlled operations should be as simple as possible, such that the state preparation step remains the main source of errors. In general, the state preparation unitaries that are feasible now are of constant depth, and the circuit size is linear in the system’s size $n$. At the first sight, it is problematic to implement the controlled reflection operation with similar resource requirements. The conventional implementation of a controlled reflection operation without the use of any ancillary qubits requires $O(n^{2})$ Toffoli and CNOT gates, while a $O(n)$ Toffoli implementation of the controlled reflection operation is possible with an additional $O(n)$ ancillary qubits [27]. Hence, a conventional implementation of controlled reflection operator is more resource-intensive than the state preparation unitary $V$, and the advantage of the CPS scheme comes at the cost of a significantly increased gate-count.
@@
Appendix D Gate Errors
@@
The cardinal task of the QSP method is to encode an estimate of the observable into the phase of the memory qubit. Therefore, we can simplify the discussion of gate errors by looking at the scenario of a faulty encoding of a phase into a single qubit. Let us assume that with probability $1-p$, $p\in[0,1]$ no error happens in the phase encoding and the true phase $\Phi_{a}$ is encoded in the single qubit. With the probability $p$ some gate errors happened, and the phase encoded in the memory qubit is an unpredictable value $\in[0,2\pi]$.
The process is repeated many times in order to achieve an estimate of $\Phi_{a}$ similar to the description following Eq. (39) in Appendix A above. However, instead of sampling from the true probability distribution corresponding to $\Phi_{a}$, we are instead sampling from the distributions
$P(X=0|{\Phi}_{N})=\frac{(1-p)}{2}(1+\cos{(2{\Phi}_{a})})+\frac{p}{2},$
$P(Y=0|{\Phi}_{N})=\frac{(1-p)}{2}(1-\sin{(2{\Phi}_{a})})+\frac{p}{2},$
where $\Phi_{a}$ is the phase encoded when no gate errors were acquired, while the second term corresponds to the case when a gate error happened, and we assume the memory qubit is left in a completely depolarised state. We can follow the procedure of the previous section to introduce (42) with the updated probabilities from Eqs. (D). All the derivations are similar in this case, and we again arrive at a variance of the estimate, which scales as $\sigma^{2}\hat{\Phi}_{a}\sim 1/M_{q}$. Note, however, that the estimated value with gate errors ($\hat{Q}_{g}$) will deviate from the true value without gate errors $\hat{Q}$ as
@@
We assume that the number of gate operations in the state preparation circuit is roughly set by the number of qubits in the $n$ (see Appendix C). The probability of a gate error happening during one QSP round is then $p\sim 3p_{1}n$, where $p_{1}$ is the single gate error probability. From this, we can estimate that the required gate error probability should be low enough such that
$\sqrt{\eta}\gtrsim\max{\{N\epsilon_{QSP},3Nn_{QSP}np_{1}\}},$ (64)
where $N\epsilon_{QSP}$ is the imprecision from the QSP approximation (see Appendix A) and $3Nn_{QSP}np_{1}$ is the inaccuracy from the gate errors. For example, to estimate an observable such as the energy of a $\text{H}_{3}^{+}$ molecule, which can be decomposed into $N=59$ Pauli strings [5] on $n=4$ qubits with a target variance of $\eta\sim 1$, the single gate error probability should be $p_{1}\lesssim 4.9\times 10^{-4}$. The amount of state preparations needed in this case is $N/\eta(\log{(N/\eta)}/\log{\log{(N/\eta)}})\sim 171$. For a bigger LiH molecule ($N=630$, $n=10$) with the same target variance, we get $p_{1}\lesssim 1.5\times 10^{-5}$, while the amount of state preparations is $\sim 2180$.
@@
For the $\text{H}_{3}^{+}$ molecule energy estimation by QEE method, the probability of single gate error should be $\lesssim 4.2\times 10^{-3}$ and the amount of state preparations is $N^{2}/\eta\sim 3481$ for a target variance of $\eta\sim 1$. For the LiH molecule the probability of the single gate error should be $\lesssim 1.5\times 10^{-4}$ and the amount of the state preparations is $3\times 10^{5}$.

### v4

@@ Appendix D Cost of implementing CPS
If we want the CPS method for observable estimation to be as generally applicable to the results of quantum algorithms, we need to satisfy the following two requirements. First requirement is that the controlled reflection and controlled Pauli operations required for $c\hat{U}_{P_{j}}$ are not more error-prone than the uncontrolled state preparation unitary $V$. Second, we need to take into account the effect of applying the state preparation unitaries sequentially for the CPS method, as opposed to a shot-noise limited multi-shot estimation scheme where the system is measured after each application of $V$. The difference between the sequential and multi-shot estimation protocols with respect to the accumulation of errors will be discussed in Appendix E below. In this section, we will focus on the implementation of controlled reflection and Pauli operations.
@@
Ideally, the implementation of the controlled operations should be as simple as possible, such that the state preparation step remains the main source of errors. The state preparation unitaries that are feasible on current platforms are of constant depth, and the circuit size is linear in the system’s size $n$, leading to an error probability that scales with $O(n)$. At the first sight, implementing a controlled reflection with error probability scaling as $O(n)$ is problematic. The conventional implementation of a controlled reflection operation without the use of any ancillary qubits requires $O(n^{2})$ Toffoli and CNOT gates, while a $O(n)$ Toffoli implementation of the controlled reflection operation is possible with an additional $O(n)$ ancillary qubits [30]. Hence, a conventional implementation of controlled reflection operator would be more resource-intensive than the state preparation unitary $V$, implying that the advantage of the CPS scheme comes at the cost of a significantly increased gate-count, compared to simply preparing the state of interest and measuring each Pauli separately.
@@
Appendix E Gate Errors
@@
The cardinal task of the QSP method is to encode an estimate of the observable into the phase of the memory qubit. Therefore, we can simplify the discussion of gate errors by looking at the scenario of a faulty encoding of a phase into a single qubit. Let us assume that with probability $1-p$, $p\in[0,1]$ no error happens in the phase encoding and the true phase $\Phi_{N}$ is encoded in the single qubit. With the probability $p$ some gate errors happened, and the phase encoded in the memory qubit is an unpredictable value $\in[0,2\pi]$.
The process is repeated many times in order to achieve an estimate of $\Phi_{a}$ similar to the description following Eq. (32) in Appendix A above. However, instead of sampling from the true probability distribution corresponding to $\Phi_{N}$, we are instead sampling from the distributions
$P(X=0|{\Phi}_{N})=\frac{(1-p)}{2}(1+\cos{(2{\Phi}_{N})})+\frac{p}{2},$
$P(Y=0|{\Phi}_{N})=\frac{(1-p)}{2}(1-\sin{(2{\Phi}_{N})})+\frac{p}{2},$
where $\Phi_{N}$ is the phase encoded when no gate errors were acquired, while the second term corresponds to the case when a gate error happened, and we assume the memory qubit is left in a completely depolarised state. We can follow the procedure of the previous section to introduce (35) with the updated probabilities from Eqs. (E). All the derivations are similar in this case, and we again arrive at a variance of the estimate, which scales as $\sigma^{2}\hat{\Phi}_{N}\sim 1/M_{q}$. Note, however, that the estimated value with gate errors ($\hat{Q}_{g}$) will deviate from the true value without gate errors $\hat{Q}$ as
@@
We assume that the errors arising in our protocol are dominated by the errors emerging in the state preparation circuits. As we state in Appendix D, we suppose that the number of gate operations in the state preparation circuit is roughly set by the number of qubits $n$. We then approximate the total probability of an error happening during the state preparation circuit as $\sim np_{1}$, where $p_{1}\ll 1$ is the average probability of making an error during a single gate operation (one-qubit or two-qubit gate). The probability of a gate error happening during one QSP round is then $p\sim 3p_{1}n$, since it involves three state preparation circuits. From this, we can estimate that the required gate error probability should be low enough such that
$\eta\gtrsim 3\sqrt{2}Nnp_{1}N_{QSP}$ (72)
holds.
@@
Similarly to the CPS, we assume that for the QPEh the probability of error is dominated by the errors of the state preparation circuit. The probability of a gate errors during QPEh for a single Pauli estimation is $\sim 3np_{1}$. Then we demand $\sqrt{\eta}\geq 3Nnp_{1}/\sqrt{\eta_{1}}$ to hold, where $\eta_{1}$ is a variance of a single Pauli estimate. Let $\eta_{1}=\eta/N$, then we get
$\eta\geq 3N^{3/2}np_{1}.$ (78)
To estimate an observable, such as the energy of molecules like $\text{H}_{3}^{+}$, LiH, $\text{H}_{2}\text{O}$, and $\text{CH}_{4}$, which can be decomposed into $N=\{59,630,1085,3887\}$ Pauli strings [6] and require $n=\{4,10,12,16\}$ qubits respectively, with a target variance of $\eta=1$, the corresponding single-gate error probabilities are summarized in Table 2.
$N$ $n$ $p_{1,\text{CPS}}$ $p_{1,\text{Heisenberg}}$ $p_{1,\text{QEE}}$
59 4 $4.49\times 10^{-5}$ $1.839\times 10^{-4}$ $4.24\times 10^{-3}$
630 10 $1.43\times 10^{-6}$ $2.11\times 10^{-6}$ $1.59\times 10^{-4}$
1085 12 $6.66\times 10^{-7}$ $7.77\times 10^{-7}$ $7.68\times 10^{-5}$
3887 16 $1.28\times 10^{-7}$ $8.50\times 10^{-8}$ $1.61\times 10^{-5}$
Table 2: The average probability of making an error during a single gate operation $p_{1}$ to achieve a fixed variance for CPS, $QPE_{h}$ and QEE methods.
The upper bounds of the amount of state preparations required for CPS method to the Heisenberg QPE given in Ref. [18] with $\alpha=1/2$, $\gamma=5/2$ constants that is stated to provide the best scaling and the QEE method are
$T_{CPS}=\frac{6\sqrt{2}N(2\pi{n}^{(1)}_{QSP}+\tilde{n}^{(1)}_{QSP})}{\sqrt{\eta}},\quad T_{QPE_{h}}=\frac{10.7N^{3/2}}{\sqrt{\eta}},\quad T_{QEE}=\frac{N^{2}}{\sqrt{\eta}}.$ (79)
The values of ${n}^{(1)}_{QSP}$ and $\tilde{n}^{(1)}_{QSP}$ are calculated numerically, solving (17) for $k$ for different values of $\epsilon_{0}$. The numerical comparison of (79) for the $\text{H}_{3}^{+}$, $LiH$, $H_{2}O$, $CH_{4}$ molecules is provided in Table 3.
Molecule $N$ $T_{{CPS}}$ $T_{QPE_{h}}$ $T_{QEE}$
$H_{3}^{+}$ 59 $1.11\times 10^{4}$ $4.85\times 10^{3}$ $3.48\times 10^{3}$
$LiH$ 630 $1.40\times 10^{5}$ $1.69\times 10^{5}$ $3.97\times 10^{5}$
$H_{2}O$ 1085 $2.50\times 10^{5}$ $3.82\times 10^{5}$ $1.177\times 10^{6}$
$CH_{4}$ 3887 $9.73\times 10^{5}$ $2.593\times 10^{6}$ $1.511\times 10^{7}$
Table 3: The amount of state preparations $T_{CPS}$, $T_{QPE_{h}}$, $T_{QEE}$ for different molecules.
One can conclude that as expected QEE is the most tolerant to the gate errors, however paying a high price of the amount of state preparations for estimating the energies of the molecules with the bigger amount of Paulis. The CPS outperforms both methods in the amount of state preparations already on the $LiH$ molecule.